OpenStudy (anonymous):

LM is tangent to circle C at point M. If CN=3x-2 NL= 4x-6 amd CL= 34 what is the length of LM? I need to know the formula or someone to explain how to work this problem. I'm lost

OpenStudy (amistre64):

these look like line measurements for a triangle but uts hard to tell with a picture to actually see.

OpenStudy (anonymous):

yes its a circle with a triangle from it. it has one side c to l and another m to l does that help

OpenStudy (amistre64):

if you can draw it in "paint" and post it in here, itd be more helpful. Paint is a basic program on many computers.

OpenStudy (amistre64):

Can you add any detail to this?

OpenStudy (anonymous):

It is a triangle with one point in the middle which is C and M is on the radius and the common point of this triangle is L.

OpenStudy (amistre64):

C is in the middle of the line? like this?

OpenStudy (anonymous):

No, the picture is of a circle hang on a minute

OpenStudy (anonymous):

its the top one

OpenStudy (amistre64):

a^2 + b^2 = c^2 (3x-2)^2 + LM^2 = [(3x-2)+(46-6)]^2 does this make sense to you?

OpenStudy (amistre64):

46-6 sposed ta be 4x-6 lol

OpenStudy (amistre64):

LM^2 = (7x-8)^2 - (3x-2)^2

OpenStudy (amistre64):

LM^2 = 49x^2 -112x +36 - 3x^2 -12x +4 LM^2 = 46x^2 -124x +40

OpenStudy (amistre64):

8*8 = 64 not 36.... you better do the math :)

OpenStudy (amistre64):

I messed up a few points there when I was expanding them.... rechack the math cause I know its wrong :)

OpenStudy (anonymous):

hang on let me liik yes but the formula is correct?

OpenStudy (amistre64):

LM^2 = 49x^2 -112x +64 -3x^2 +12x -4 the formula is correct, that I know.

OpenStudy (anonymous):

thank you sooo much I will be needing more help later, Im having problems finding this in my math book

OpenStudy (amistre64):

the pic you sent, the radius is (3x-2) so just use it as the base, then add the other 2 lengths together to get the hypotenuse, and find the missing leg...

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

im going to work this out and get back with u

OpenStudy (amistre64):

ok..

OpenStudy (anonymous):

how did u get 112x

OpenStudy (amistre64):

when you add the lengths together for the hypotenuse you get: (3x-2) + (4x-6) 3x -2 +4x -6 = 7x -8 (7x-8)^2 = 49x^2 - 56(2)x +64 49^2 -112x +64

OpenStudy (anonymous):

let me check

OpenStudy (anonymous):

oh I see where now

OpenStudy (anonymous):

I think Im still lost... I believe I am doing the x^2 wrong

OpenStudy (amistre64):

doing the x^2 wrong.... now your sounding as crazy as me...

OpenStudy (anonymous):

Im afraid to ask

OpenStudy (amistre64):

so what you have there is a simple triangle that has a length of 2 sides and you want to find the third side.... one side has been given to you in 2 seperate equation, and the shortest leg is simply the length of the radius which is one of the equations they gave you for the other.... plug it into the pythagoreum thrm and get your answer :)

OpenStudy (anonymous):

ok let me do it

OpenStudy (anonymous):

I believe the final answer is 16. but I will have to recheck the process again. I have to leave for now. thanks for all the help I'll be back again.